OpenStudy (anonymous):

The shampoo you like to use is on sale. But you only get the lower price if you buy 3 bottles...? ...You want to know how much you'll save What is the formula? s=sales price per bottle r=regular price per bottle c=cost savings

OpenStudy (apoorvk):

c=3r-3s 3, since you buy 3 bottles. c is the total cost savings.

OpenStudy (anonymous):

it will be c= 3(r-s)

hero (hero):

The best way to go about this is to make up a scenario: Suppose shampoo costs $5 per bottle, which would amount to $15 if you purchased thee bottles. However, you can buy 3 bottles for $12 To find the sales price, you would divide $12/3 = $4. So you're basically saving $3 so in essence r = $5 s = $4 c = $3 ($1 per bottle) As for a formula: It will in general be Regular Price - Sales Price = Cost Savings...so r - s = c In this case $5 - $4 = $1 meaning you'll save $1 per bottle.

OpenStudy (apoorvk):

nah. but you need to buy 3 bottles remember? @hero

hero (hero):

apoorvk, are you sure you read everything I posted? If you're saving $1 per bottle, and you buy three bottles, $1 times 3 = a total cost savings of $3. I already addressed buying 3 bottles. If you have a better explanation, you should provide one rather than trying to negate mine.

hero (hero):

Okay fine... For three bottles: 3(r - s) = 3c

OpenStudy (apoorvk):

ohkay. i somehow missed the last line. sorry, my bad!

OpenStudy (apoorvk):

but, 'c' is the total savings already, and not the savings per bottle, so i guess just "c=3(r-s)" shall do.

hero (hero):

@apoorvk , you speak too fast r = $5 s = $4 c = $1 3(r - s) = 3c 3(5 - 4) = 3 15 - 12 = 3 3 = 3

hero (hero):

c = $1 in this case I originally posted c = $3 The correct equation stands at 3(r - s) = 3c

OpenStudy (apoorvk):

umm. right, i did speak too fast there. dumb me. apologies once again. silly me!